Suppose we create a 99% confidence interval for a sample size of 20 units with mean 200 standard deviation 30, what are the upper and lower limits of the interval?

A> 186.85 and 213.15
B> 184.37 and 215.63
C> 184.24 and 215.76
D> 182.69 and 217.30
E> None of the above
1
Expert's answer
2020-03-20T11:41:07-0400

We need to construct the 99% confidence interval for the population mean "\mu" with known population standard deviation "\sigma."

Given that "\bar{X}=200, \sigma=30, n=20, \alpha=0.01."

The critical value for "\alpha=0.01" is "z_c=z_{1-\alpha/2}=2.576\approx2.58"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\cdot{\sigma \over \sqrt{n}}, \bar{X}+z_c\cdot{\sigma \over \sqrt{n}})\approx"

"\approx(200-2.58\cdot{30 \over \sqrt{20}}, 200+2.58\cdot{30 \over \sqrt{20}})\approx"

"\approx(182.69,\ 217.30)"

D) "182.69" and "217.30"


We need to construct the 99% confidence interval for the population mean "\mu" with unknown population standard deviation "(\sigma)" for which reason the sample standard deviation "(s)" is used instead.

Given that "\bar{X}=200,s=30, n=20, \alpha=0.01."

The critical value for "\alpha=0.01, df=20-1=19" is "t_c=2.861."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-t_c\cdot{s \over \sqrt{n}}, \bar{X}+t_c\cdot{s \over \sqrt{n}})\approx"

"\approx(200-2.861\cdot{30 \over \sqrt{20}}, 200+2.861\cdot{30 \over \sqrt{20}})\approx"

"\approx(180.81,\ 219.19)"

E) None of the above